Joe can clean his room in 84 minutes what is joes work rate?
Answer by psbhowmick(878)

84 mins = 84/60 = 7/5 hours

Joe can clean the room in 7/5 hours.
So in 1 hour Joe can clean 5/7 of the room.

Joe's work rate is 5/7 per hour.